OpenStudy (anonymous):

how do you solve proportions using cross products? can you maybe give me an example aswell?

OpenStudy (anonymous):

When the terms of a proportion are cross multiplied, the cross products are equal. Cross multiplication is the multiplication of the numerator of the first ratio by the denominator of the second ratio and the multiplication of the denominator of the first ratio by the numerator of the second ratio. 21/3=3/10 21*10 = 7-*3 201= 210 i think this is what u were asking for?